Yvette is making a circular tablecloth with diameter of 5 feet.To the nearest square foot,how much cloth will be used in the tablecloth?
Answer by checkley79(3341)

AREA=PI*R^2
AREA=3.14*5^2
AREA=3.14*25
AREA=78.5 FT^2.